The smelting department of Company X has the following production and cost data for September:

• Production: Beginning work in process 200 units that are 100% complete as to materials and 82% complete as to conversion costs; units started into production 1,000 units; units transferred out to next department 950 units; ending work in process are 100% complete as to materials and 42% complete as to conversion costs.

• Manufacturing costs: Beginning work in process TL8,200 materials, TL9,270 labor, and TL2,200 manufacturing overhead costs. In the current period, TL50,500 worth of materials, TL25,700 worth of labor and TL32,000 worth of manufacturing overhead costs are incurred.

•Question: Compute the material cost portion of the transferred out inventory if the company uses FIFO method to compute equivalent units.
Answer #1

Equivalent units = (950-200) + 250 x 100% = 1000 units

Cost per unit = 50500 / 1000 = 50.5 per unit

Material Cost portion of units transferred = 8200 + 750 x 50.5 = 46075